Use the markup equation S = C + rC, where S is the selling price, C is the cost, and r is the markup rate.
A pair of jeans with a selling price of $40 has a markup rate of 60%. Find the cost of the pair of jeans.

Answer by mananth(12270)

S = C + rC,
s=40
r=60%
40 = C+0.6c
40=1.6c
40/1.6=c
c= 25 cost
m.ananth@hotmail.ca